A jogger completes a circular track with a diameter of 325 m in 12 minutes. What is the average velocity of the jogger?
Answer #1

verage speed = Distance
Answer #2

Average speed = Distance
Answer #3

Average velocity will be zero because displacement will be zero for a complete revolution.
Answer #4

avg velocity=s/t=pi d/t=325x3.14/12x60=1.42m/s
Answer #5

the average velocity of the jogger is 0,06 m/s
Answer #6

v=50m/s
Answer #7

the average velocity of the jogger v=299 m/s
Answer #8

Average speed = Distance
Answer #9

the average velocity of the jogger is 1,104 m/s